# A Composite-Loss Graph Neural Network for the Multivariate Post-Processing of Ensemble Weather Forecasts

### v1

@@ 5.4 Multivariate performance of visibility forecasts
Figure 12 shows the mean ES and the corresponding skill scores of raw, POLR, and multivariate post-processed visibility forecasts relative to GNN-ES. As illustrated in Figure 12a, all post-processed forecasts improve upon the raw ensemble predictions, although this advantage diminishes with increasing lead time. In order to highlight differences between post-processing models, in Figure 12b the GNN-ES is used as the reference. Relative to this benchmark, dualGNN shows only a 0.4% improvement, and it is the only model with a positive skill score; nevertheless, the performance of dualGNN and GNN-ES can be considered effectively equivalent. The strongest POLR variant, POLR-ECC, lags slightly behind the reference by 2%, closely followed by POLR-SSh, representing a further 0.2% decrease in the mean skill score. The weakest post-processing model is POLR, which exhibits an average 9.94% decrease relative to GNN-ES.
A slightly different pattern observed in Figure 13, which displays the mean VS of raw, POLR, and multivariate post-processed visibility forecasts, along with the corresponding skill scores, again using GNN-ES as the reference. Unlike in Figure 12a, clear improvements over the raw ensemble are now observed only for the GNN models that minimize multivariate scores and the two-step POLR methods. As before, this advantage diminishes with increasing lead time. Based on the mean VS, the models are ranked as follows: dualGNN, POLR-SSh, POLR-ECC, GNN-ES, POLR, GNN-CRPS, and finally the raw ensemble.
@@ 6 Discussion
For both weather variables, the composite-loss GNN (dualGNN) consistently improved upon all reference models, achieving the best performance according to multivariate scores while maintaining well-calibrated marginal distributions. In contrast, empirical copula-based methods did not yield a single clear winner, whereas dualGNN achieved the best performance across both ES and VS. In the case of visibility forecasts, dualGNN also achieved the lowest mean CRPS, surpassed only by a GNN specifically optimized for CRPS. Furthermore, the rank structure of forecasts generated by dualGNN captured relevant dependencies among stations, producing the most uniform rank histograms. Being nonparametric, the method is applicable to any weather variable and allows flexibility in the number of ensemble members generated.
@@
Acknowledgments.  This work was supported by the EKÖP-25-4-II University Research Scholarship Program of the Ministry for Culture and Innovation, funded by the National Research, Development and Innovation Fund.

### v2

@@ 5.4 Multivariate performance of visibility forecasts
Figure 12 shows the mean ES and the corresponding skill scores of POLR, and multivariate post-processed visibility forecasts relative to GNN-ES, with the raw forecasts omitted from the skill score plot to better highlight the differences between the post-processed models. As illustrated in Figure 12a, all post-processed forecasts improve upon the raw ensemble predictions, although this advantage diminishes with increasing lead time. In order to highlight differences between post-processing models, in Figure 12b the GNN-ES is used as the reference. Relative to this benchmark, dualGNN shows only a 0.4% improvement, and it is the only model with a positive skill score; nevertheless, the performance of dualGNN and GNN-ES can be considered effectively equivalent. The strongest POLR variant, POLR-ECC, lags slightly behind the reference by 2%, closely followed by POLR-SSh, representing a further 0.2% decrease in the mean skill score. The weakest post-processing model is POLR, which exhibits an average 9.94% decrease relative to GNN-ES.
A slightly different pattern observed in Figure 13, which displays the mean VS of POLR, and multivariate post-processed visibility forecasts, along with the corresponding skill scores, again using GNN-ES as the reference, with the raw forecasts omitted from Figure 13b. Unlike in Figure 12a, clear improvements over the raw ensemble are now observed only for the GNN models that minimize multivariate scores and the two-step POLR methods. As before, this advantage diminishes with increasing lead time. Based on the mean VS, the models are ranked as follows: dualGNN, POLR-SSh, POLR-ECC, GNN-ES, POLR, GNN-CRPS, and finally the raw ensemble.
@@ 6 Discussion
For both weather variables, the composite-loss GNN (dualGNN) consistently improved upon all reference models, achieving the best performance according to multivariate scores while maintaining well-calibrated marginal distributions. In contrast, empirical copula-based methods did not yield a single clear winner, whereas dualGNN achieved the best performance across both ES and VS. In the case of visibility forecasts, dualGNN also achieved the lowest mean CRPS, surpassed only by a GNN specifically optimized for CRPS. Furthermore, the rank structure of forecasts generated by dualGNN captured relevant dependencies among stations, producing the most uniform rank histograms. Being nonparametric, the method is applicable to any weather variable and allows flexibility in the number of ensemble members generated. Overall, the dualGNN consistently generated accurate forecasts across target variables with diverse characteristics. This demonstrates that the model can robustly handle variables with markedly different statistical properties while sustaining high predictive skill and preserving meaningful rank dependencies.
@@
Acknowledgments.  The author gratefully acknowledges the support of the National Research, Development, and Innovation Office under Grant No. K142849. Furthermore, she thanks her former supervisor, Sándor Baran, for kindly reviewing and providing valuable feedback on the manuscript.